Newman Weatherspoon, a super salesman contemplating retirement on his fifty-fifth birthday, decides to create a fund on a 9% basis that will enable him to withdraw $18,380 per year on June 30, beginning in 2024 and continuing through 2027. To develop this fund, Newman intends to make equal contributions on June 30 of each of the years 2020–2023.

How much must the balance of the fund equal on June 30, 2023, in order for Newman to satisfy his objective?

What are each of Newman’s contributions to the fund?

Note: I am using a financial calculator to solve this. How would this question be solved using this calculator?
Answer #1

(A)

Annual withdrawal from from = 2024 to 2027 = $18380

Number of withdrawal= 4

Annual interest rate= 9%

Fund balance of June 30 2023= present value of annual withdrawal.

=. Uniform withdrawal* (1-(1+r)-n)/r

=18380*(1-1.09-4)/0.09

=18380*3.23972(see annuity table)= 59546

B)

Number of contribution = 4

Future value of contribution= 59546

Future value of annuity of 1= annual contribution*((1+r)n​​​​​-1)/r

=59546=annual contribution*(1.094​​​​​-1)/0.09

Annual contribution = 59546/4.573129= 13021